Use the integrating factor to solve the differential equation. dz/dy = z tan y + sin y
1
Expert's answer
2021-09-07T02:40:52-0400

"\cfrac{dz}{dy}=z \cdot tan(y)+sin(y)
\\ \cfrac{dz}{dy}-z \cdot tan(y)=sin(y)"


Then, when we compare the form to "\cfrac{dz}{dy}+P \cdot z=Q", we can identify that the integrating factor will be: "I=e^{\int Pdy}=e^{-\int tan(y)dy}=e^{\ln(cos(y))}=cos(y)"


Then, we multiply the integrating factor I and proceed to solve the equation:

"\\ (cos(y))[\cfrac{dz}{dy}-z \cdot tan(y)]=sin(y)cos(y)
\\ cos(y)\cfrac{dz}{dy}-z \cdot sin(y)=sin(y)cos(y)
\\ \cfrac{d}{dy}(z \cdot cos(y))=sin(y)cos(y) \iff d(z \cdot cos(y))= sin(y)cos(y){dy}
\\ \int d(z \cdot cos(y))=\int sin(y)cos(y){dy}
\\ z \cdot cos(y)=-\frac{1}{2} cos^2(y)+C"


We divide the last result by cos(y) to determine z:

"z =C\cdot sec(y)-\cfrac{1}{2} cos(y)"


In conclusion, we determine that "z =C\cdot sec(y)-\cfrac{cos(y)}{2}" .